OpenStudy (anonymous):

how many pounds are there in 1 Newton?

OpenStudy (anonymous):

N=Kg/m^2

OpenStudy (anonymous):

The pound or pound-mass (abbreviations: lb, lbm, lbm, ℔[1]) is a unit of mass used in the imperial, United States customary and other systems of measurement. A number of different definitions have been used, the most common today being the international avoirdupois pound which is legally defined as exactly 0.45359237 kilograms.

OpenStudy (noelgreco):

It's Googleable.

OpenStudy (shane_b):

You can also just calculate it:\[1kg = 2.20462lbs\]\[1N=\frac{2.20462lbs}{9.80655m/s^2}=0.22481...lb_f\]

OpenStudy (anonymous):

1kg=2.20462lbs ...

OpenStudy (unklerhaukus):

\[1[\text N]=0.22481[\text{lbf}]\]